what is the function for y=-3*2x;x=3

Found 2 solutions by Alan3354, Fombitz:
Answer by Alan3354(69443)

what is the function for y=-3*2x;x=3
----
y = -3*2x --> y = -6x
Not clear what you mean.

Answer by Fombitz(32388)